# PoSh: Using Scene Graphs to Guide LLMs-as-a-Judge for Detailed Image Descriptions

### v1

@@ Abstract
While vision-language models (VLMs) have advanced into detailed image description, evaluation remains a challenge. Standard metrics (e.g. CIDEr, SPICE) were designed for short texts and tuned to recognize errors that are now uncommon, such as object misidentification. In contrast, long texts require sensitivity to attribute and relation attachments and scores that localize errors to particular text spans. In this work, we introduce PoSh, a metric for detailed image description that uses scene graphs as structured rubrics to guide LLMs-as-a-Judge, producing aggregate scores grounded in fine-grained errors (e.g. mistakes in compositional understanding). PoSh is replicable, interpretable and a better proxy for human raters than existing metrics (including GPT4o-as-a-Judge). To validate PoSh, we introduce a challenging new dataset, DOCENT. This novel benchmark contains artwork, paired with expert-written references, and model-generated descriptions, augmented with granular and coarse judgments of their quality from art history students. Thus, DOCENT enables evaluating both detailed image description metrics and detailed image description itself in a challenging new domain. We show that PoSh achieves stronger correlations (+$0.05$ Spearman $\rho$) with the human judgments in DOCENT than the best open-weight alternatives, is robust to image type (using CapArena, an existing dataset of web imagery) and is a capable reward function, outperforming standard supervised fine-tuning. Then, using PoSh, we characterize the performance of open and closed models in describing the paintings, sketches and statues in DOCENT and find that foundation models struggle to achieve full, error-free coverage of images with rich scene dynamics, establishing a demanding new task to gauge VLM progress. Through both PoSh and DOCENT, we hope to enable advances in important areas such as assistive text generation. We make our metric and our benchmark available at https://github.com/amith-ananthram/posh.
@@ 1 Introduction
A picture is worth a thousand words – can vision-language models (VLMs) capture all of them? VLMs have saturated traditional image understanding benchmarks from short captioning to question answering (Li et al., 2025). New, more challenging tasks are needed to measure VLM progress. Detailed image description is of particular interest as it requires comprehensive understanding – e.g., in Fig. 1, a VLM must correctly specify who is pouring the water. This deep perception is a better proxy for the demands of the real world, where diverse user queries may not be reflected in VQA benchmarks (Chen et al., 2024). Moreover, it enables meaningful applications such as image assistive (“alt”) text generation that could greatly expand accessibility online (Mack et al., 2021).
However, making progress on detailed description requires cheap, reliable methods for scoring models. Human evaluation is costly, involving the painstaking comparison of long texts. Even so, there is often no substitute as most metrics were designed for short texts and older models (Berger et al., 2024). Moreover, while metrics that produce a single coarse score of overall quality allow for the ranking of models, they offer little insight into the granular issues driving performance. Granular issues include mistakes in each generation, like the positions of the people in Fig. 1, and omissions in each reference, like the details of the bird’s beak in Fig. 3. Automatically localizing such errors is critical as long generations with similar coarse scores may differ in multiple dimensions of interest (e.g., facial features, body orientations, etc.). Otherwise, prompt and/or model iteration necessitates expensive manual inspection to understand which description aspects need improvement.
@@
In this work, we propose PoSh  [^1] PoSh (PrOofing Scene grapHs) can judge if your detailed descriptions are what you (really really) want., a metric for evaluating detailed descriptions that addresses these challenges. PoSh extracts scene graphs from a generated description and its reference to use as structured rubrics for an LLM to granularly identify mistakes and omissions (see Fig. 2), pinpointing the textual spans containing errors like attribute/relation mis-attachment. Then, it aggregates these localized errors into coarse scores for mistakes, omissions and overall quality. Thus, PoSh weds the strengths of structured methods like scene graphs (Anderson et al., 2016), which reduce descriptions to their consequential visual components, with the strengths of LLMs/VLMs-as-a-Judge (Zheng et al., 2023), which flexibly compare these visual components against diverse surface realizations.
@@
As PoSh’s coarse scores are grounded in its granular scores, it is interpretable, providing clear insights into the errors driving model performance. Moreover, because PoSh is entirely open-weight, it is inexpensive to use and replicable, an important pre-requisite for both adoption by researchers and deployment by practitioners that is not afforded by closed models.
@@
Efforts to introduce metrics for longer generations have been constrained by a lack of human judgments, especially at a granular scale and for diverse imagery (see Table 1). To address this, we introduce DOCENT, a novel benchmark whose focus is visual art. DOCENT contains paintings, sketches and sculptures with expert-written assistive text that exhaustively describes features like clothing, physical orientation, relative positioning and gaze, drawn from the U.S. National Gallery of Art (see Figs. 2 and 3). It includes generations from current VLMs with judgments from art history students of their mistakes, omissions and overall quality at two resolutions: granular and coarse. Thus, DOCENT enables evaluating description metrics and descriptions themselves [^2] AI research often uses caption and alt-text interchangeably. However, according to Web Content Accessibility Guidelines, captions are related to an image while alt-text conveys the information in an image. As our focus is evaluating generations that could serve as alt-text, we use the term description..
@@
We validate PoSh against the human judgments in DOCENT. We show that PoSh recovers human description rankings more often (+$3$ percentage points) and achieves stronger correlations with human-derived scores (+$0.05$ Spearman $\rho$) than existing overlap and open-weight alternatives (e.g. SPICE, CAPTURE, LLaVa-Critic), even surpassing GPT4o-as-a-Judge. Moreover, using judgments in CapArena (Cheng et al., 2025), we show this strength is robust to image type. Then, given its calibration, we experiment with using PoSh as a reward function for describing the images in DOCENT and find that this yields meaningfully better descriptions than supervised fine-tuning (SFT).
@@
Finally, using PoSh, we characterize the performance of open and closed models in describing the artwork in DOCENT, establishing a difficult new task. In so doing, we extend detailed description to a technically challenging and socially impactful domain: assistive text generation for artwork, whose visual complexity and diversity stress VLMs (Bengamra et al., 2024) (see Fig. 1).
@@ 2 Related Work
Image description is under-specified – the correct way to describe an image is often task-specific. This is especially true for assistive text which has context-dependent requirements (Kreiss et al., 2022). Moreover, in such sensitive applications, correlated failures between reference-free metrics and VLMs relying on similar components could prove dangerous to end users (Deutsch et al., 2022). Thus, our focus is reference-based evaluation. Traditional metrics were not designed to evaluate long text and can involve truncation due to limited context length (e.g. CLIPScore) (Papineni et al., 2002; Lin, 2004; Banerjee & Lavie, 2005; Vedantam et al., 2015; See et al., 2017; Hessel et al., 2021; Sarto et al., 2023). Recent work has explored LLMs/VLMs-as-Judges though this requires potentially expensive API calls and offers limited replicability (Chan et al., 2023; Cheng et al., 2025). Even when replicable, they do not provide interpretable, grounded granular scores (Xiong et al., 2025).
While prior metrics like SPICE and CAPTURE leverage scene graphs, they forgo their rich structure by ignoring object attachment (Anderson et al., 2016; Dong et al., 2024). This favors generations with misattributed details (as in Fig. 1). In summarization, Scialom et al. (2021) use question generation and answering (QA) to compare a summary and its source. In text-to-image generation, Cho et al. (2024) use GPT4 to extract and verify a scene graph from a visual prompt. PoSh extends these approaches to detailed description evaluation that is replicable and interpretable. With small models, it extracts scene graphs to use as structured rubrics for guiding an open-weight LLM-as-a-Judge.
@@
Evaluating such a metric requires human judgments of model generations. Though there are many detailed image description benchmarks (Urbanek et al., 2024; Onoe et al., 2024; Garg et al., 2024; Lu et al., 2025; Ye et al., 2025), most release no such judgments. One notable exception is CapArena which contains coarse rankings of descriptions for web imagery (Cheng et al., 2025). In contrast, our new dataset, DOCENT contains both granular and coarse judgments, enabling the evaluation of fine-grained metrics like PoSh. Moreover, it expands detailed description to artwork whose scene dynamics and expert-written references are considerably more complex (see Table 1).
@@ 3 PoSh: A New Metric for Detailed Image Description
As in SPICE (Anderson et al., 2016), given a description $d$, a scene graph $G(d)$ is a structured representation of $d$. Specifically, $G(d)=\left<O(d),E(d),K(d)\right$> where $O(d)\subseteq C$ is a set of objects, $E(d)\subseteq O(d)\times A$ is a set of attributes associated with each object and $K(d)\subseteq O(d)\times R\times O(d)$ are a set of relation edges between objects. $C$, AA and $R$ are open-world sets of all possible object, attribute and relation classes.
Given a generation gen with its reference ref we extract sentence-level scene graphs $G_{i}(\textit{gen}),G_{j}(\textit{ref})$ for each using off-the-shelf dependency parsing and combine them via coreference resolution (Honnibal et al., 2020; Martinelli et al., 2024). This produces scene graphs with full coverage of each text where each component is localized to text spans, allowing for grounded, interpretable scoring. We provide pseudocode for this extraction in Section A.1.1.
@@
We note this is a natural place to introduce tunable weights (as in Dong et al. (2024)) to adapt PoSh to particular datasets. As we aim to demonstrate robustness, we leave these terms unweighted.
@@ 4 DOCENT: A New Benchmark for Detailed Description of Art
DOCENT is a benchmark for evaluating detailed description metrics and detailed descriptions themselves. It consists of $1,750$ works of art with expert-written references from the Open Data Program at the U.S. National Gallery of Art (NGA) [^4] https://www.nga.gov/open-access-images/open-data.html. For $100$ of these images, we produce four generations from current small and frontier VLMs and collect $300$ granular (for $75$ images) and $600$ coarse judgments from annotators knowledgeable in art of mistakes and omissions [^5] We forgo fluency as recommended by Kasai et al. (2022). On average, coarse judgments took $5$ minutes and granular judgments took $18$ minutes (six annotation days). This highlights both the cost of manual evaluation and the need for metrics that are reliable proxies.
@@
The accompanying references are detailed descriptions whose purpose is accessibility – as such, they follow guidelines that include tips for describing color (e.g., “color can be likened to temperature”) and handling ambiguity (e.g, “describe what makes something ambiguous”). These context informed requirements highlight the need for reference based metrics (Kreiss et al., 2022)666www.nga.gov/visit/accessibility/collection-image-descriptions.
Compared to existing detailed image description benchmarks, DOCENT contains considerably more visual complexity (see Table 1). On average, its images contain $16\%$ more objects and nearly twice as many people [^7] As measured by OneFormer Jain et al. (2023) who require description of their orientation, features, clothing, etc. Consequently, the average length and scene graph size of its reference descriptions are nearly double.
@@
We generate detailed descriptions for $100$ images in DOCENT from four current VLMs that span transparency and model size (from open data/open weight to frontier models): LLaVA-1.6-7B (Liu et al., 2024), Molmo-D-7B (Deitke et al., 2024), GPT4o and Claude Sonnet 3.5. A metric that discriminates among these generations similarly to their human judgments could gauge progress in detailed image description in small and large VLMs over time. Additional details (prompts, date of API access) can be found in Section A.2.1.
@@
Given the complexity of our images and the detail of their expert descriptions, we recruit $24$ art history undergraduate majors, masters students and PhD students with domain familiarity to provide high quality judgments of generations. All annotators were sighted with full color vision and native speakers of English. They were compensated at a rate of \$2222/hour for their time. [^8] This study was conducted under an approved IRB which will be specified upon publication.
@@
Half of our annotators identify mistakes and omissions in our model generations. For each image, an annotator is shown its reference and then its four model generations in random order. First, they look at the image, read the reference and then the current generation. Next, by selecting narrow text spans, annotators first identify mistakes in the generation (i.e. precision errors) and then omissions in the reference that are not in the generation (i.e. recall errors). When identifying omissions, as in Kasai et al. (2022), we ask annotators to mentally correct narrow mistakes in the generation first to avoid double-penalizing a model for both incorrect specificity and lack of specificity. For example, a generation that describes a woman as a man is an error in precision but not in recall. We include our task instructions and interface [^9] Hosted on Label Studio (https://labelstud.io) screenshots in Figs. 4 and 6.
@@
The other half of our annotators provide coarse judgments of our model generations. For a given image, an annotator is shown its reference and two generations (#1 and #2) in random order and asked to rank the generations in terms of mistakes (i.e. precision), omissions (i.e. recall) and overall quality. These pairwise judgments avoid some of the inter-annotator inconsistency of Likert ratings, especially for long text (Novikova et al., 2018).
@@
For our granular judgments, as recommended by Hripcsak & Rothschild (2005) for span annotation tasks where the boundaries of negative examples (i.e. non-errors) are ill-defined, we measure agreement using the relaxed F1 (matching spans that contain 50% overlapping tokens). Under this measure, our student annotators exhibit strong agreement among themselves and with our expert.
Our coarse judgments exhibit moderate inter-annotator agreement, with Krippendorf’s $\alpha=0.509$, $0.409$ and $0.459$ for mistakes, omissions and overall quality (Landis & Koch, 1977). This level of agreement is unsurprising for coarse detailed description evaluation – judgment requires weighing the relative importance of each text’s granular errors and is consequently more subjective. Nevertheless, our student annotators exhibit moderate to strong correlations with our expert, with significant Pearson $\rho$ values of $0.727$, $0.501$ and $0.492$ for mistakes, omissions and overall quality respectively.
@@
When considering the performance of the four models included in DOCENT, we observe expected trends, adding to our confidence in the quality of our judgments: the smaller models make more mistakes and have more omissions than the larger models (see Tables 5 and 5). Though most models make few mistakes, they all struggle with omissions. The best model, gpt4o covers only $44.0\%$ of the visual information conveyed in DOCENT’s references. Raising this requires continued prompt iteration, highlighting the need for an automated metric that can reliably measure both granular and coarse differences in mistakes and omissions.

### v2

@@ Abstract
While vision-language models (VLMs) have advanced into detailed image description, evaluation remains a challenge. Standard metrics (e.g. CIDEr, SPICE) were designed for short texts and tuned to recognize errors that are now uncommon, such as object misidentification. In contrast, long texts require sensitivity to attribute and relation attachments and scores that localize errors to particular text spans. In this work, we introduce PoSh, a metric for detailed image description that uses scene graphs as structured rubrics to guide LLMs-as-a-Judge, producing aggregate scores grounded in fine-grained errors (e.g. mistakes in compositional understanding). PoSh is replicable, interpretable and a better proxy for human raters than existing metrics (including GPT4o-as-a-Judge). To validate PoSh, we introduce a new dataset, DOCENT. This novel benchmark contains artwork, paired with expert-written references, and model-generated descriptions, augmented with granular and coarse judgments of their quality from art history students. Thus, DOCENT enables evaluating both detailed image description metrics and detailed image description itself in a challenging new domain. We show that PoSh achieves stronger correlations (+$0.05$ Spearman $\rho$) with the human judgments in DOCENT than the best open-weight alternatives, is robust to image type (using CapArena, an existing dataset of web imagery) and is a capable reward function, outperforming standard supervised fine-tuning. Then, using PoSh, we characterize the performance of open and closed models in describing the paintings, sketches and statues in DOCENT and find that foundation models struggle to achieve full, error-free coverage of images with rich scene dynamics, establishing a demanding new task to gauge VLM progress. Through both PoSh and DOCENT, we hope to enable advances in important areas such as assistive text generation. We make our metric and our benchmark available at https://github.com/amith-ananthram/posh.
@@ 1 Introduction
A picture is worth a thousand words – can vision-language models (VLMs) capture all of them? VLMs have saturated traditional image understanding benchmarks from short captioning to question answering (li2025benchmark). New, more challenging tasks are needed to measure VLM progress. Detailed image description is of particular interest as it requires comprehensive understanding – e.g., in Fig. 1, a VLM must correctly specify who is pouring the water. This deep perception is a better proxy for the demands of the real world, where diverse user queries may not be reflected in VQA benchmarks (chenwe). Moreover, it enables meaningful applications such as image assistive (“alt”) text generation that could greatly expand accessibility online (mack2021we).
However, making progress on detailed description requires cheap, reliable methods for scoring models. Human evaluation is costly, involving the painstaking comparison of long texts. Even so, there is often no substitute as most metrics were designed for short texts and older models (berger2024surveying). Moreover, while metrics that produce a single coarse score of overall quality allow for the ranking of models, they offer little insight into the granular issues driving performance. Granular issues include mistakes in each generation, like the positions of the people in Fig. 1, and omissions in each reference, like the details of the bird’s beak in Fig. 3. Automatically localizing such errors is critical as long generations with similar coarse scores may differ in multiple dimensions of interest (e.g., facial features, body orientations, etc.). Otherwise, prompt and/or model iteration necessitates expensive manual inspection to understand which description aspects need improvement.
@@
In this work, we propose PoSh  [^1] PoSh (PrOofing Scene grapHs) can judge if your detailed descriptions are what you (really really) want., a metric for evaluating detailed descriptions that addresses these challenges. PoSh extracts scene graphs from a generated description and its reference to use as structured rubrics for an LLM to granularly identify mistakes and omissions (see Fig. 2), pinpointing the textual spans containing errors like attribute/relation mis-attachment. Then, it aggregates these localized errors into coarse scores for mistakes, omissions and overall quality. Thus, PoSh weds the strengths of structured methods like scene graphs (anderson2016spice), which reduce descriptions to their consequential visual components, with the strengths of LLMs/VLMs-as-a-Judge (zheng2023judging), which flexibly compare these visual components against diverse surface realizations.
@@
As PoSh’s coarse scores are grounded in its granular scores, it is interpretable, providing clear insights into the errors driving model performance. Moreover, because PoSh is entirely open-weight, it is inexpensive to use and perfectly replicable, an important pre-requisite for both adoption by researchers and deployment by practitioners that is not afforded by closed models.
@@
Efforts to introduce metrics for longer generations have been constrained by a lack of human judgments, especially at a granular scale and for diverse imagery (see Table 1). To address this, we introduce DOCENT, a novel benchmark whose focus is visual art. DOCENT contains paintings, sketches and sculptures with expert-written assistive text that exhaustively describes features like clothing, physical orientation, relative positioning and gaze, drawn from the U.S. National Gallery of Art (see Figs. 2 and 3). It includes generations from current VLMs with judgments from art history students of their mistakes, omissions and overall quality at two resolutions: granular and coarse. Thus, DOCENT enables evaluating description [^2] AI research often uses caption and alt-text interchangeably. However, according to Web Content Accessibility Guidelines, captions are related to an image while alt-text conveys the information in an image. As our focus is evaluating generations that could serve as alt-text, we use the term description. metrics and descriptions themselves.
@@
We validate PoSh against the human judgments in DOCENT. We show that PoSh recovers human description rankings more often (+$3$ percentage points) and achieves stronger correlations with human-derived scores (+$0.05$ Spearman $\rho$) than existing overlap and open-weight alternatives (e.g. SPICE, CAPTURE, LLaVa-Critic), even surpassing GPT4o-as-a-Judge. Moreover, using judgments in CapArena (cheng2025caparena), we show this strength is robust to image type. Then, given its calibration, we experiment with using PoSh as a reward function for describing the images in DOCENT and find that this yields meaningfully better descriptions than supervised fine-tuning (SFT).
@@
Finally, using PoSh, we characterize the performance of open and closed models in describing the artwork in DOCENT, establishing a difficult new task. In so doing, we extend detailed description to a technically challenging and socially impactful domain: assistive text generation for artwork, whose visual complexity and diversity stress VLMs (bengamra2024comprehensive) (see Fig. 1).
@@ 2 Related Work
Image description is under-specified – the correct way to describe an image is often task-specific. This is especially true for assistive text which has context-dependent requirements (kreiss2022context). Moreover, in such sensitive applications, correlated failures between reference-free metrics and VLMs relying on similar components could prove dangerous to end users (deutsch2022limitations). Thus, our focus is reference-based evaluation. Traditional metrics were not designed to evaluate long text and can involve truncation due to limited context length (e.g. CLIPScore) (papineni2002bleu; lin2004rouge; banerjee2005meteor; vedantam2015cider; see2017get; hessel2021clipscore; sarto2023positive). Recent work has explored LLMs/VLMs-as-Judges though this requires potentially expensive API calls and offers limited replicability (chan2023clair; cheng2025caparena). Even when replicable, they do not provide interpretable, grounded granular scores (xiong2024llavacritic).
While prior metrics like SPICE and CAPTURE leverage scene graphs, they forgo their rich structure by ignoring object attachment (anderson2016spice; dong2024benchmarking). This favors generations with misattributed details (as in Fig. 1). In summarization, scialom2021questeval use question generation and answering (QA) to compare a summary and its source. In text-to-image generation, chodavidsonian use GPT4 to extract and verify a scene graph from a visual prompt. PoSh extends these approaches to detailed description evaluation that is replicable and interpretable. With small models, it extracts scene graphs to use as structured rubrics for guiding an open-weight LLM-as-a-Judge.
@@
Evaluating such a metric requires human judgments of model generations. Though there are many detailed image description benchmarks (urbanek2024picture; onoe2024docci; garg-etal-2024-imageinwords; lu2025benchmarking; ye2025painting), most release no such judgments. One notable exception is CapArena which contains coarse rankings of descriptions for web imagery (cheng2025caparena). In contrast, our new dataset, DOCENT contains both granular and coarse judgments, enabling the evaluation of fine-grained metrics like PoSh. Moreover, it expands detailed description to artwork whose scene dynamics and expert-written references are considerably more complex (see Table 1).
@@ 3 PoSh: A New Metric for Detailed Image Description
As in SPICE (anderson2016spice), given a description $d$, a scene graph $G(d)$ is a structured representation of $d$. Specifically, $G(d)=\left<O(d),E(d),K(d)\right$> where $O(d)\subseteq C$ is a set of objects, $E(d)\subseteq O(d)\times A$ is a set of attributes associated with each object and $K(d)\subseteq O(d)\times R\times O(d)$ are a set of relation edges between objects. $C$, AA and $R$ are open-world sets of all possible object, attribute and relation classes.
Given a generation gen with its reference ref we extract sentence-level scene graphs $G_{i}(\textit{gen}),G_{j}(\textit{ref})$ for each using off-the-shelf dependency parsing and combine them via coreference resolution (spacy; martinelli-etal-2024-maverick). This produces scene graphs with full coverage of each text where each component is localized to text spans, allowing for grounded, interpretable scoring. We provide pseudocode for this extraction in Section A.1.2.
@@
We note this is a natural place to introduce tunable weights (as in dong2024benchmarking) to adapt PoSh to particular datasets. As we aim to demonstrate robustness, we leave these terms unweighted.
@@ 4 DOCENT: A New Benchmark for Detailed Description of Art
DOCENT is a benchmark for evaluating detailed description metrics and detailed descriptions themselves. It consists of $1,750$ works of art with expert-written references from the Open Data Program at the U.S. National Gallery of Art (NGA) [^4] https://www.nga.gov/open-access-images/open-data.html. For $100$ of these images, we produce four generations from current small and frontier VLMs and collect $300$ granular (for $75$ images) and $600$ coarse judgments from annotators knowledgeable in art of mistakes and omissions [^5] We forgo fluency as recommended by kasai-etal-2022-transparent. On average, coarse judgments took $5$ minutes and granular judgments took $18$ minutes (six annotation days). This highlights both the cost of manual evaluation and the need for metrics that are reliable proxies.
@@
The accompanying references are detailed descriptions whose purpose is accessibility – as such, they follow guidelines666www.nga.gov/visit/accessibility/collection-image-descriptions that include tips for describing color (e.g., “color can be likened to temperature”) and handling ambiguity (e.g, “describe what makes something ambiguous”). These context informed requirements highlight the need for reference based metrics (kreiss2022context).
Compared to existing detailed image description benchmarks, DOCENT contains considerably more visual complexity (see Table 1). On average, its images contain $16\%$ more objects and nearly twice as many people [^7] As measured by OneFormer jain2023oneformer who require description of their orientation, features, clothing, etc. Consequently, the average length and scene graph size of its reference descriptions are nearly double.
@@
We generate detailed descriptions for $100$ images in DOCENT from four current VLMs that span transparency and model size (from open data/open weight to frontier models): LLaVA-1.6-7B (liu2024improved), Molmo-D-7B (deitke2024molmo), GPT4o and Claude Sonnet 3.5. A metric that discriminates among these generations similarly to their human judgments could gauge progress in detailed image description in small and large VLMs over time. Additional details (prompts, date of API access) can be found in Section A.2.1.
@@
Given the complexity of our images and the detail of their expert descriptions, we recruit $24$ art history undergraduate majors, masters students and PhD students with domain familiarity to provide high quality judgments of generations. All annotators were sighted with full color vision and native speakers of English. They were compensated at a rate of \$2222/hour for their time. [^8] This study was conducted under Columbia University IRB protocol AAAV6216.
@@
Half of our annotators identify mistakes and omissions in our model generations. For each image, an annotator is shown its reference and then its four model generations in random order. First, they look at the image, read the reference and then the current generation. Next, by selecting narrow text spans, annotators first identify mistakes in the generation (i.e. precision errors) and then omissions in the reference that are not in the generation (i.e. recall errors). When identifying omissions, as in kasai-etal-2022-transparent, we ask annotators to mentally correct narrow mistakes in the generation first to avoid double-penalizing a model for both incorrect specificity and lack of specificity. For example, a generation that describes a woman as a man is an error in precision but not in recall. We include our task instructions and interface [^9] Hosted on Label Studio (https://labelstud.io) screenshots in Figs. 4 and 6.
@@
The other half of our annotators provide coarse judgments of our model generations. For a given image, an annotator is shown its reference and two generations (#1 and #2) in random order and asked to rank the generations in terms of mistakes (i.e. precision), omissions (i.e. recall) and overall quality. These pairwise judgments avoid some of the inter-annotator inconsistency of Likert ratings, especially for long text (novikova2018rankme).
@@
For our granular judgments, as recommended by hripcsak2005agreement for span annotation tasks where the boundaries of negative examples (i.e. non-errors) are ill-defined, we measure agreement using the relaxed F1 (matching spans that contain 50% overlapping tokens). Under this measure, our student annotators exhibit strong agreement among themselves and with our expert.
Our coarse judgments exhibit moderate inter-annotator agreement, with Krippendorf’s $\alpha=0.509$, $0.409$ and $0.459$ for mistakes, omissions and overall quality (landis1977measurement). This level of agreement is unsurprising for coarse detailed description evaluation – judgment requires weighing the relative importance of each text’s granular errors and is consequently more subjective. Nevertheless, our student annotators exhibit moderate to strong correlations with our expert, with significant Pearson $\rho$ values of $0.727$, $0.501$ and $0.492$ for mistakes, omissions and overall quality respectively.
@@
When considering the performance of the four models included in DOCENT, we observe expected trends, adding to our confidence in the quality of our judgments: the smaller models make more mistakes and have more omissions than the larger models (see Tables 5 and 5). Though most models make few mistakes, they all struggle with omissions. The best model, gpt4o covers only $50.1\%$ of the visual information conveyed in DOCENT’s references. Raising this requires continued prompt iteration, highlighting the need for an automated metric that can reliably measure both granular and coarse differences in mistakes and omissions.